OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the graphing method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (2 points) Part 3: Explain what the solution represents on the graph. (1 point) 3x + 6y = 18 2x + 5y = 20

OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the graphing method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (2 points) Part 3: Explain what the solution represents on the graph. (1 point) 3x + 6y = 18 2x + 5y = 20

OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. (4 points) Part 2: Provide the solution to the system. (2 points) 5x – 3y = 4 3x – 5y = –4

OpenStudy (anonymous):

I'm not good at English, so I don't know how to express it in perfect way. However, I know how to solve it by graphing. first of all, for the first equation, you make a table to figure out the points on the Cartesian coordinate. to my experience, it's a line. from that line, for any values of x , you have a value of y, for example 

OpenStudy (anonymous):

do you know what I mean? by graphing, you can find out any y from the graph, which corresponds to particular x

OpenStudy (anonymous):

can yo do this one''

OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. (4 points) Part 2: Provide the solution to the system. (2 points) 5x – 3y = 4 3x – 5y = –4

OpenStudy (anonymous):

the table is something like